OpenStudy (anonymous):

sgadi, what is the solution of this equation.. The product of two positive numbers is 16. Find the numbers if the sum is least.

OpenStudy (sgadi):

the answer is 4,4 ..... but I am not able to get the proof

OpenStudy (anonymous):

Thanks for the hwlkp.. haha

OpenStudy (sgadi):

welcome

OpenStudy (anonymous):

ok so you have two numbers x, y x+ y = O BUT xy=16 where O is some number, and we want it to be a min so you want to find an equation with one variable since we know xy=16 so solve for x we get x = 16/y then plug into the unknown we get 16/y+y = O take the derivative and set equal to zero we get -16/y^2 + 1 = 0 solving for y you get y=+4, -4. since we are dealing with positive numbers, take y = 4. but then we need x. so you plug back into xy=16 4*x = 16 therefore x =4

OpenStudy (sgadi):

thanks :-)

OpenStudy (anonymous):

yeah i forgot how to give a "proof"

OpenStudy (anonymous):

dude if you dont understand i can explain again if you need me to